OpenStudy (anonymous):

what does this mean? find two consecutive positive even numbers the sum of whose squares is 100.

OpenStudy (anonymous):

so it's looking for two numbers (that are positive) and even, so 2, 4, 6, 8, 10. etc. and the squares of those numbers will add up to 100 so the square would be the numbers multiplied by themselves. they have to be smaller than 10, since 10 * 10 itself = 100

OpenStudy (anonymous):

so then we can go up, so they'll be in ORDER and their even, so there's the options of; 2 and 4, 4 and 6, 6 and 8 the squares of 2 and 4 only add up to 20 so that's toolittle let's try 4 and 6 16 + 36 = 52, so too little 6 and 8 36 + 64 = 100 so that's the answer :) 6 and 8

OpenStudy (amistre64):

good trial and error approach :) another approach is to define 2 even consecutive numbers as: 2n and 2n+2 (2n)^2 + (2n+2)^2 = 100 4n^2 + 4n^2 +4 + 8n = 100 8n^2 + 8n + 4 = 100 8n^2 + 8n - 96 = 0 n^2 + n - 12 = 0 n = -4 or 3, -4 is out so let n=3 and we get the results already posted :)